OpenStudy (anonymous):

HELP!!!! Can someone please help me with this?? Line L is perpendicular to a line m with equation -6x + 2y = -48. Line L intersects line m at (7, -3). The equation of line L, in standard form is, ________???

OpenStudy (anonymous):

you gotta find the slope of \(-6x+2y=-48\) first before you can find the slope of the perpendicular line do you know it?

OpenStudy (anonymous):

Yes, I know that part

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

for the slope i mean

OpenStudy (anonymous):

1/3x -y = -.75 I'm not sure if it's right

OpenStudy (anonymous):

For standard form

OpenStudy (anonymous):

no it is not right

OpenStudy (anonymous):

oops no it isn't

OpenStudy (anonymous):

the slope of \(-6x+2y=-48\) is \(3\)

OpenStudy (anonymous):

Ok, I got that part, but I need the line perpendicular to that

OpenStudy (anonymous):

your line, being perpendicular, will have a slope of \(-\frac{1}{3}\) the "negative reciprocal"

OpenStudy (anonymous):

and you have to write it in "standard form" correct?

OpenStudy (anonymous):

What is the equation of the perpendicular line in Y-intercept form? Yes, in standard form

OpenStudy (anonymous):

start with the point slope form \[y=y_1=m(x-x_1)\] using \(x_1=7,y_1=-3,m=-\frac{1}{3}\)

OpenStudy (anonymous):

I already have the slope

OpenStudy (anonymous):

yeah and the point too

OpenStudy (anonymous):

What do you mean by the point?

OpenStudy (anonymous):

Line L intersects line m at (7, -3)

OpenStudy (anonymous):

so the point is \((7,-3)\)

OpenStudy (anonymous):

ok, I got that

OpenStudy (anonymous):

you want to start with \[y+3=-\frac{1}{3}(x-7)\]

OpenStudy (anonymous):

How do you get from there to the standard form?

OpenStudy (anonymous):

i would multiply everything by 3 to get rid of obnoxious fraction

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I don't know how to do that.

OpenStudy (anonymous):

\[3(y+3)=3\times -\frac{1}{3}(x-7)\]

OpenStudy (anonymous):

then the fraction is gone, you get \[3y+9=-(x-7)\] now distribute the minus sign on the right

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

No, we have not learned this way

OpenStudy (anonymous):

not sure what you learned, but this is what you need if you are to put in in standard form

OpenStudy (anonymous):

Ok 3y+9=-x+7

OpenStudy (anonymous):

We have weekly quizzes, but are not taught in class for them

OpenStudy (anonymous):

ok then to put the x on the left, add x to both sides \[x+3y+9=7\] and finally subtract 9

OpenStudy (anonymous):

I don't know how to do that

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

x+3y=-2

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

btw i don't believe you do not know how to subtract 9 just saying ..

OpenStudy (anonymous):

I am exhausted right now, sorry

OpenStudy (anonymous):

So, what do I do from there? Is that the final answer?

OpenStudy (anonymous):

now you have a snack, you are done

OpenStudy (anonymous):

WHEW!!! Thanks!!!!!! You are awesome!!!